. Assume all the values are in 8-bit 2s complement representation. Perform the following operations

. i. 01111111 + 11111111

ii. 00100101 - 10110111
Answer #1

ans i :

(01111111)2 = (127)10

(11111111)2 = (-1)10 (Taking 2s complement and attaching a negative sign.)

(127)10 + (-1)10 = 127 - 1 = (126)10

(126)10 = (01111110)2

So the answer is 01111110.

ans ii :

(00100101)2 = (37)10

(10110111)2 = (-73)10 (Taking 2s complement and attaching a negative sign.)

(37)10 - (-73)10 = 37 - (-73) = 37 + 73 = (110)10

(110)10 = (01101110)2

So the answer is 01101110.